# Emergent morphogenesis via planar fabrication enabled by a reduced model of composites

### v1

@@ Abstract
pacs:
Valid PACS appear here
I Introduction
Rapid and cost-effective manufacturing of 3D structures from planar 2D materials is challenging and the manufacturing methods are in significant demand in fields such as soft robotics, automobile manufacturing, biomedical devices, aerospace components, and deployable structures. The existing structures fabricated from active and responsive materials have enabled applications in aerospace furuya1992concept ; hanaor2001evaluation ; santiago2013advances , medicine du2018programmable ; banerjee2020origami ; babaee2019temperature , and intelligent machines wang2017modular ; chi2022bistable , but they remain primarily material-focused. There is a growing need for scalable manufacturing frameworks that can accelerate real-world deployment. In this context, computational design – particularly approaches that enable rapid simulation and optimization – plays a pivotal role. Our work addresses this challenge by introducing a reduced computational model that enables efficient, accurate simulation of shape-morphing composites, leveraging algorithms originally developed in the computer graphics community grinspun2003discrete .
A variety of pathways have been explored to drive shape morphing, including tailored material designs dudek2025shape , composite structures daynes2013review ; rodriguez2016shape , microstructurally engineered polymers aharoni2018universal ; jin2022shape ; studart2014bioinspired , origami-inspired mechanisms hawkes2010programmable ; tolley2014self , and architected metastructures boston2022spanwise ; risso2022highly ; zhang2025shape . The transformation from 2D to 3D forms is governed by the interplay between stretching and bending armon2011geometry , where strain mismatches embedded in planar preforms can lead to a variety of 3D shapes aharoni2018universal ; ware2015voxelated ; giudici2022multiple ; modes2011gaussian . In particular, bilayer composites comprising a stimuli-responsive material bonded to a patterned kirigami layer mungekar2024directed offer a scalable,
manufacturing-driven strategy for programmable shape morphing across a broad range of applications, such as deployable structures, patient-specific biomedical devices, and adaptive surfaces. In this work, we use a bilayer system composed of a Shrinky Dink thermoplastic and an inert kirigami-patterned plastic as a model platform for programmable morphable surfaces. Our reduced-order numerical method, demonstrated on this system, can be readily augmented to simulate and guide the design of other classes of shape-morphing composite shells.
@@ I Introduction
In this paper, we introduce a numerical framework for efficient and accurate prediction of stimulus-driven 3D shape morphing in bilayer composite shells. Unlike conventional approaches that require explicit modeling of multiple layers, we represent the entire layered composite using a single layer of nodes and formulate an effective energy model that captures the coupled mechanics of the bilayer system. Our method builds on the discrete shells approach grinspun2003discrete , where the total elastic energy of a thin structure is computed from in-plane stretching (changes in edge lengths) and out-of-plane bending (dihedral angles between adjacent triangular elements), each referenced to rest lengths and rest angles that define the undeformed, stress-free configuration. We advance this framework by allowing these rest lengths and rest angles to evolve explicitly as functions of the applied stimulus (heat in this work), enabling the model to capture permanent, stimulus-driven deformations such as those caused by thermal contraction in heat-shrinking layers. This energy-based approach yields accurate predictions of complex 3D configurations arising from 2D patterned precursors, without resorting to computationally expensive full 3D simulations. We validate our framework by comparing numerical predictions to physical experiments on bilayer composites composed of a heat-shrinking thermoplastic (Shrinky Dink) and a layer of 3D printed PLA (polylactic acid).
Figure 1 shows representative 3D shapes fabricated in physical experiments (left) and their numerically simulated counterparts (right). The experimental fabrication protocol builds directly on prior work mungekar2024directed , while the present study provides, for the first time, a predictive numerical framework for these systems. Our main contributions are:
@@
II Results and Discussion
II.1 Emergent 3D Shapes via Planar Fabrication of Bilayer Composites
@@
Figure 3 illustrates the complete fabrication process. First, the kirigami pattern is designed based on the desired final shape; for example, Figure 3A shows a six-arm star on a circular thermoplastic sheet. The pattern is then digitally modeled with the prescribed thickness using Fusion fusion2025 and fabricated by 3D printing, as shown in Figure 3B. Next, the printed PLA pattern is bonded to a precisely cut thermoplastic substrate, and the assembled bilayer composite is cured for over 10 hours to ensure robust adhesion (Figure 3C). After curing, thermal actuation is used to induce shape morphing, as shown in Figure 3D, by heating the planar composite in an oven for approximately 15 seconds. The response of the Shrinky Dink layer is characterized by measuring its shrinkage as a function of normalized temperature (Figure 3E), where the temperature is scaled by the glass transition temperature ($T_{g}$) of polystyrene, an amorphous thermoplastic without a sharp melting point. The sheet begins to shrink when $T/T_{g}>1$ (temperatures are measured in Kelvins), but the shrink ratio $L/L_{0}$ does not stabilize until it exceeds 1.15, corresponding to temperatures above $300^{\circ}$F, as indicated by a vertical dashed line in Figure 3E. Consequently, all experiments are performed at temperatures greater than $300^{\circ}$F.
@@
II.2 Numerical Simulation with a Reduced Model
Motivated by the experimental observations of programmable shape morphing in bilayer composites – comprising a 3D printed PLA kirigami layer bonded to a Shrinky Dink sheet – we propose a new energy formulation that couples planar stretching with out-of-plane deformation. We first briefly review the discrete elastic shells method in Section II.2.1, and then introduce our reduced energy model for composite shells in Section II.2.2.
II.2.1 Discrete Elastic Shells method
@@
In the discrete shells framework, the total elastic energy is the sum of stretching energies (associated with each edge) and bending energies (associated with each hinge) savin_growth_2011 :
@@
The strains are fully determined by the DOF vector $\mathbf{X}$ and are invariant with respect to the global coordinate system. Specifically, the axial strain for each edge is $\varepsilon_{i}=l_{i}/l_{0,i}-1$, where $l_{i}$ is the current edge length and $l_{0,i}$ is its rest (reference) length in the undeformed mesh. The dihedral angle $\theta_{j}$ is defined as the angle between the normals of the two triangles sharing hinge $j$. The stretching and bending stiffnesses are
@@
where $Y$ is the Young’s modulus, $h$ is the shell thickness, and $l_{0}$ is the mean edge length. These discrete expressions are constructed such that, in the limit of mesh refinement, the model recovers the continuum Föppl–von Kármán equations for nonlinear thin plates. The three composites have domains $\Omega$ of 100 mm radius (pattern A), 100 mm $\times$ 156 mm (pattern B), and 100 mm $\times$ 100 mm (pattern C), Figures 2 and 4. With the mesh configurations in Figure 4, we have $l_{0}=3.2$ $m$, $4.0$ mm, and $5.8$ mm for A, B, and C, respectively. The thicknesses of two layers and their corresponding effective stretching and bending stiffness values, $k_{s}$ and $k_{b}$, are given in Section II.3.
@@
In our model, the computational mesh consists of two distinct types of regions, Figure 4 A-C(i): (i) bilayer regions $\mathcal{B}$, where both the Shrinky Dink and the inert kirigami layer are present, and (ii) single-layer regions $\Omega\setminus\mathcal{B}$, where only the Shrinky Dink substrate remains (corresponding to the cut-out areas of the kirigami pattern). The material properties – such as Young’s modulus $Y$, and consequently the stretching and bending stiffnesses $k_{s}$ and $k_{b}$ – are assigned locally to each mesh element based on its region: bilayer regions use effective properties of the combined layers, while single-layer regions use properties of the Shrinky Dink alone. The specific numerical values used for all material parameters are provided in the Section II.3.
@@
While the assignment of material properties enables accurate representation of spatial heterogeneity in the composite, classical discrete plate models are limited to flat, stress-free reference states. In contrast, discrete shells extend this framework by permitting nonzero natural (rest) curvature, achieved by introducing nonzero rest dihedral angles. This is implemented by replacing the bending term $\frac{1}{2}\,k_{b}\,\theta_{j}^{2}$ with $\frac{1}{2}\,k_{b}\,\left(\theta_{j}-\theta_{j}^{\mathrm{rest}}\right)^{2}$ in the energy expression. Such an extension is essential for capturing the permanent, stimulus-induced shape changes observed in our bilayer composites. Building on these principles, we next present our reduced energy model for programmable composite shells, which incorporates stimulus-dependent rest configurations to predict the full range of morphable shapes observed in experiments.
@@
Figure 4: Reference-configurations meshes for designs A-C in Figures 1 and 2. A-C(i) Thermoplastic PLA bilayer boundaries (solid lines) on discrete mesh. A-C(ii) The distribution of thermal strains as defined in Equation 7. See Section II.2.2 for definitions of $\Omega$, $\mathcal{B}$, $\varepsilon_{pre}^{\mathrm{th}}$, and $d_{i}$. D. Local node and edge IDs for a bending pair. The number of nodes $N_{n}$, number of edges $N_{e}$, and number of triangular elements $N_{t}$, $(N_{n},N_{e},N_{t})$ are (970, 2800, 1831) for pattern A, (1215, 3505, 2291) for pattern B, and (388, 1087, 700) for pattern C, respectively.
II.2.2 Energy Formulation for Stimulus-Responsive Composite Shells
@@
The core innovation of our approach is to extend the discrete shells framework so that the elastic energy depends explicitly on both the DOF vector $\mathbf{X}$ – representing the shape of the structure – and the applied external stimulus, such as temperature. This enables us to efficiently model the permanent, stimulus-driven morphing of bilayer shells using only a single layer of mesh nodes, without the need for computationally intensive 3D elements. The key physical mechanism is the non-uniform thermal strain induced in the bilayer composite: when a 3D-printed PLA pattern is bonded to a Shrinky Dink substrate, only the Shrinky Dink contracts under heating, while the PLA remains largely inert. This mismatch in planar contraction leads to locally varying natural curvatures – captured in our model as stimulus-dependent rest dihedral angles at the mesh hinges. By properly designing the kirigami pattern, we control which regions bend and by how much, enabling programmable shape transformation. Building on the framework established above, we generalize the energy expression in Eq. 4 by allowing the rest dihedral angle $\bar{\theta}_{j}$ to depend explicitly on the applied thermal stimulus, $T$.
@@
where $\varepsilon_{i}^{\mathrm{th}}(T)$ is the stimulus-induced thermal strain in the Shrinky Dink layer (measured from experiments, cf. Figure 3E). The first term accounts for in-plane stretching, while the second term models out-of-plane bending. Here, $\theta_{j}(\bm{X})$ is the dihedral angle at hinge $j$, and $\beta l_{0}\Delta\varepsilon_{j}(\bm{X},T)$ replaces the rest dihedral angle $\bar{\theta}_{j}$ in Eq. 4, quantifying the local mismatch in planar strain across the two triangles sharing each hinge. The parameter $\beta$ controls the strength of this coupling and is determined by fitting to experimental data (see Supporting Information for details). Specifically, $\beta$ (unit $m^{-1}$) denotes the coupling effect per unit triangle‐side length, and $\beta l_{0}$ gives the coupling strength for a mesh whose triangles have a mean side length $l_{0}$. For convenience, we set $\beta=1/l_{0}$ here. With this definition, we aim to diminish the dependence of $\beta$ on numerical discretization. The explicit expression for $\Delta\varepsilon_{j}$ is provided later in this section. We write $\varepsilon_{i}(\bm{X})$, $\theta_{j}(\bm{X})$, and $\Delta\varepsilon_{j}(\bm{X},T)$ to emphasize that these quantities are computed from the DOF vector $\bm{X}$ and, where relevant, the applied stimulus $T$. Similarly, $\varepsilon_{i}^{\mathrm{th}}(T)$ denotes the thermal strain as a function of $T$, computed directly from the experimentally measured contraction of the Shrinky Dink sheet at each temperature (see Figure 3E).
The thermal strain results from a mismatch caused by shrinkage due to temperature change, is described by $\varepsilon_{pre}^{\mathrm{th}}=\alpha\Delta T$. By observing experimental results, we found that deformation around the boundary of the Shrinky Dink sheet is more significant than in the region occupied by the PLA composite. Another contributing factor is the glue effect. In some areas, the glue has spread beyond the PLA layer, making the composite stiffer than the pure Shrinky Dink sheet. Consequently, the magnitude of the thermal shrinkage strain increases with the distance from the PLA region, where it remains zero. This variation of thermal strain is given by

### v2

@@ Abstract
keywords:
planar fabrication, bilayer composite, thermal actuation, reduced model, experimental validation
Yupeng Zhang Adam Alon Mohammad Khalid Jawed*
[…]
{affiliations}
Yupeng Zhang, Adam Alon, Mohammad Khalid Jawed
Department of Mechanical & Aerospace Engineering
University of California Los Angeles
Los Angeles, CA 90095, USA
Email Address: khalidjm@seas.ucla.edu
11footnotetext: These authors contributed equally to this work.
1 Introduction
Rapid and cost-effective manufacturing of 3D structures from planar 2D materials is challenging and the manufacturing methods are in significant demand in fields such as soft robotics, automobile manufacturing, biomedical devices, aerospace components, and deployable structures. The existing structures fabricated from active and responsive materials have enabled applications in aerospace [21, 24, 42], medicine [19, 8, 7], and intelligent machines [50, 14], but they remain primarily material-focused. There is a growing need for scalable manufacturing frameworks that can accelerate real-world deployment. In this context, computational design – particularly approaches that enable rapid simulation and optimization – plays a pivotal role. Our work addresses this challenge by introducing a reduced computational model that enables efficient, accurate simulation of shape-morphing composites, leveraging algorithms originally developed in the computer graphics community [23].
A variety of pathways have been explored to achieve planar to three dimensional shape morphing, including soft kirigami composites for flexible deployable structures [53], GAN-based inverse design [11, 12], shape optimization of electrostatic zipper actuator [3], magnetic kirigami dome metasheet for dynamic shape shifting [13], multimodal shape evolution of shape memory polymers [52], tailored material designs [20], composite structures [18, 40], microstructurally engineered polymers [2, 27, 44], origami-inspired mechanisms [25, 48], and architected metastructures [9, 39, 54]. In particular, Brzin and Brojan [11] developed a mechanical analog framework for the inverse design of soft morphing composite beams based on strain mismatch induced by temperature actuation. The transformation from 2D to 3D forms is governed by the interplay between stretching and bending [4], where strain mismatches embedded in planar preforms can lead to a variety of 3D shapes [2, 51, 22, 34]. In particular, bilayer composites comprising a stimuli-responsive material bonded to a patterned kirigami layer [35] offer a scalable, manufacturing driven strategy for programmable shape morphing across a broad range of applications, such as deployable structures, patient-specific biomedical devices, and adaptive surfaces. In this work, we use a bilayer system composed of a Shrinky Dink thermoplastic and an inert kirigami-patterned plastic as a model platform for programmable morphable surfaces. Our reduced-order numerical method, demonstrated on this system, can be readily augmented to simulate and guide the design of other classes of shape-morphing composite shells.
@@ 1 Introduction
In this paper, we introduce a numerical framework for efficient and accurate prediction of stimulus-driven 3D shape morphing in bilayer composite shells. Unlike conventional approaches that require explicit modeling of multiple layers, we represent the entire layered composite using a single layer of nodes and formulate an effective energy model that captures the coupled mechanics of the bilayer system. Our method builds on the discrete shells approach [23], where the total elastic energy of a thin structure is computed from in-plane stretching (changes in edge lengths) and out-of-plane bending (dihedral angles between adjacent triangular elements), each referenced to rest lengths and rest angles that define the undeformed, stress-free configuration. We advance this framework by allowing these rest lengths and rest angles to evolve explicitly as functions of the applied stimulus (heat in this work), enabling the model to capture permanent, stimulus-driven deformations such as those caused by thermal contraction in heat-shrinking layers. This energy-based approach enables accurate predictions of complex 3D configurations arising from 2D patterned precursors, without resorting to computationally expensive full 3D simulations. A direct comparison of computational cost between the discrete elastic shell (DES) and finite element (FE) methods is presented at the end of Section 2.5. We validate our framework by comparing numerical predictions to physical experiments on bilayer composites composed of a heat-shrinking thermoplastic (Shrinky Dink) and a layer of 3D printed PLA (polylactic acid).
Figure 1 shows representative 3D shapes fabricated in physical experiments (left) and their numerically simulated counterparts (right). The experimental fabrication protocol builds directly on prior work [35], while the present study provides, for the first time, a predictive numerical framework for these systems. Our main contributions are:
@@
2 Results and Discussion
2.1 Emergent 3D Shapes via Planar Fabrication of Bilayer Composites
@@
Figure 3 illustrates the complete fabrication process. First, the kirigami pattern is designed based on the desired final shape; for example, Figure 3A shows a six-arm star on a circular thermoplastic sheet. The pattern is then digitally modeled with the prescribed thickness using Fusion [6] and fabricated by 3D printing, as shown in Figure 3B. Next, the printed PLA pattern is bonded to a precisely cut thermoplastic substrate, and the assembled bilayer composite is cured for over 10 hours to ensure robust adhesion (Figure 3C). After curing, thermal actuation is used to induce shape morphing, as shown in Figure 3D, by heating the planar composite in an oven for approximately 15 seconds. The response of the Shrinky Dink layer is characterized by measuring its shrinkage as a function of normalized temperature (Figure 3E), where the temperature is scaled by the glass transition temperature ($T_{g}$) of polystyrene, an amorphous thermoplastic without a sharp melting point. The sheet begins to shrink when $T/T_{g}>1$ (temperatures are measured in Kelvins), but the shrink ratio $L/L_{0}$ does not stabilize until it exceeds 1.15, corresponding to temperatures above $300^{\circ}$F, as indicated by a vertical dashed line in Figure 3E. Consequently, all experiments are performed at temperatures greater than $300^{\circ}$F.
@@
2.2 Numerical Simulation with a Reduced Model
Motivated by the experimental observations of programmable shape morphing in bilayer composites – comprising a 3D printed PLA kirigami layer bonded to a Shrinky Dink sheet – we propose a new energy formulation that couples planar stretching with out-of-plane deformation. We first briefly review the discrete elastic shells method in Section 2.2.1, and then introduce our reduced energy model for composite shells in Section 2.2.2.
2.2.1 Discrete Elastic Shells Method
@@
In the discrete shells framework, the total elastic energy is the sum of stretching energies (associated with each edge) and bending energies (associated with each hinge) [43]:
@@
The strains are fully determined by the DOF vector $\mathbf{X}$ and are invariant with respect to the global coordinate system. Specifically, the axial strain for each edge is $\varepsilon_{i}=l_{i}/l_{0,i}-1$, where $l_{i}$ is the current edge length and $l_{0,i}$ is its rest (reference) length in the undeformed mesh. The dihedral angle $\theta_{j}$ is defined as the angle between the normals of the two triangles sharing hinge $j$. The stretching stiffness $k_{s}$ and bending stiffness $k_{b}$ are
@@
where $Y$ is the effective Young’s modulus, $h$ is the shell thickness, and $l_{0}$ is the mean edge length. These discrete expressions are constructed such that, in the limit of mesh refinement, the model recovers the continuum Föppl–von Kármán equations for nonlinear thin plates. The three composites have domains $\Omega$ of 100 mm radius (pattern A), 100 mm $\times$ 156 mm (pattern B), and 100 mm $\times$ 100 mm (pattern C), Figures 2 and 4. With the mesh configurations in Figure 4, we have $l_{0}=3.2$ $m$, $4.0$ mm, and $5.8$ mm for A, B, and C, respectively. The thicknesses of two layers and their corresponding effective stretching and bending stiffness values, $k_{s}$ and $k_{b}$, are given in Section 2.3.
@@
In our model, the computational mesh consists of two distinct types of regions, Figure 4 A-C(i): (i) bilayer regions $\mathcal{B}$, where both the Shrinky Dink and the inert kirigami layer are present, and (ii) single-layer regions $\Omega\setminus\mathcal{B}$, where only the Shrinky Dink substrate remains (corresponding to the cut-out areas of the kirigami pattern). The material properties – such as Young’s modulus $Y$, and consequently the stretching and bending stiffnesses $k_{s}$ and $k_{b}$ – are assigned locally to each mesh element based on its region: bilayer regions use effective properties of the combined layers, while single-layer regions use properties of the Shrinky Dink alone. The specific numerical values used for all material parameters are provided in the Section 2.3.
@@
While the assignment of material properties enables accurate representation of spatial heterogeneity in the composite, classical discrete plate models are limited to flat, stress-free reference states. In contrast, discrete shells extend this framework by permitting nonzero natural (rest) curvature, achieved by introducing nonzero rest dihedral angles. This is implemented by replacing the bending term $\frac{1}{2}\,k_{b}\,\theta_{j}^{2}$ with $\frac{1}{2}\,k_{b}\,\left(\theta_{j}-\bar{\theta}_{j}\right)^{2}$ in the energy expression. Such an extension is essential for capturing the permanent, stimulus-induced shape changes observed in our bilayer composites. Building on these principles, we next present our reduced energy model for programmable composite shells, which incorporates stimulus-dependent rest configurations to predict the full range of morphable shapes observed in experiments.
@@
Figure 4: Reference-configurations meshes for designs A-C in Figures 1 and 2. A-C(i) Thermoplastic PLA bilayer boundaries (solid lines) on discrete mesh. A-C(ii) The distribution of thermal strains as defined in Equation 16. See Section 2.2.2 for definitions of $\Omega$, $\mathcal{B}$, $\varepsilon_{\rm pre}^{\mathrm{th}}$, and $d_{i}$. D. Local node and edge IDs for a bending pair. The number of nodes $N_{n}$, number of edges $N_{e}$, and number of triangular elements $N_{t}$, $(N_{n},N_{e},N_{t})$ are (970, 2800, 1831) for pattern A, (1215, 3505, 2291) for pattern B, and (388, 1087, 700) for pattern C, respectively. E. Variation of normalized thermal strain $\varepsilon^{\mathrm{\rm th}}/\varepsilon_{\rm pre}^{\mathrm{\rm th}}$ with normalized distance ${d}/{d_{\max}}$ for different values of $\gamma$, corresponding to the characteristic length $L_{c}=d_{\max}/\gamma$ in Equation 12. F(i)-(iii) Deformed configurations of pattern A with $\varepsilon_{\rm pre}^{\mathrm{\rm th}}=0.7$ for $\gamma=0.1$, 1, and $5$, respectively. G. Top views of the corresponding configurations shown in F.
2.2.2 Energy Formulation for Stimulus-Responsive Composite Shells
@@
The core innovation of our approach is to extend the discrete shells framework so that the elastic energy depends explicitly on both the DOF vector $\mathbf{X}$ – representing the shape of the structure – and the applied external stimulus, such as temperature. This enables us to efficiently model the permanent, stimulus-driven morphing of bilayer shells using only a single layer of mesh nodes, without the need for computationally intensive 3D elements. The key physical mechanism is the non-uniform thermal strain induced in the bilayer composite: when a 3D-printed PLA pattern is bonded to a Shrinky Dink substrate, only the Shrinky Dink contracts under heating, while the PLA remains largely inert. This mismatch in planar contraction leads to locally varying natural curvatures – captured in our model as stimulus-dependent rest dihedral angles at the mesh hinges. By properly designing the kirigami pattern, we control which regions bend and by how much, enabling programmable shape transformation. Building on the framework established above, we generalize the energy expression in Equation 4 by allowing the rest dihedral angle $\bar{\theta}_{j}$ to depend explicitly on the applied thermal stimulus, $T$.
@@
where $\varepsilon_{i}^{\mathrm{th}}(T)$ is the stimulus-induced thermal strain in the Shrinky Dink layer (measured from experiments, cf. Figure 3E). The first term accounts for in-plane stretching, while the second term models out-of-plane bending. Here, $\theta_{j}(\bm{X})$ is the dihedral angle at hinge $j$, and $\beta l_{0}\Delta\varepsilon_{j}(\bm{X},T)$ replaces the rest dihedral angle $\bar{\theta}_{j}$ in Equation 4, quantifying the local mismatch in planar strain across the two triangles sharing each hinge. The explicit expression for $\Delta\varepsilon_{j}$ is provided later in this section. We write $\varepsilon_{i}(\bm{X})$, $\theta_{j}(\bm{X})$, and $\Delta\varepsilon_{j}(\bm{X},T)$ to emphasize that these quantities are computed from the DOF vector $\bm{X}$ and, where relevant, the applied stimulus $T$. Similarly, $\varepsilon_{i}^{\mathrm{th}}(T)$ denotes the thermal strain as a function of $T$, computed directly from the experimentally measured contraction of the Shrinky Dink sheet at each temperature (see Figure 3E).
The parameter $\beta$ quantifies the strength of coupling between in-plane strain mismatch and out-of-plane bending and has units of inverse length. We consider a bilayer slender beam with traction-free boundary conditions subjected to uniform tensile strains $\varepsilon_{1}$ and $\varepsilon_{2}$ in layer 1 and layer 2, respectively. When the strain mismatch $\Delta\varepsilon=\varepsilon_{1}-\varepsilon_{2}$ is zero, the beam remains flat, i.e. corresponds to a zero natural curvature. Whereas a nonzero mismatch $\Delta\varepsilon\neq 0$ induces a nonzero natural curvature $\bar{\kappa}\neq 0$. In the discrete shell formulation, the curvature $\bar{\kappa}$ is related to the bending angle $\bar{\theta}$ between adjacent elements by [23]
$\bar{\kappa}=\bar{\theta}/l_{0},$ (7)
where $l_{0}$ is the element length. From the energy in Equation 6, we have $\bar{\theta}=\beta l_{0}\Delta\varepsilon$. Substitute this into Equation 7, we obtain
$\bar{\kappa}=\beta\Delta\varepsilon,$ (8)
indicating that $\beta$ directly relates local strain mismatch to local curvature per unit length.
In principle, $\beta$ could be calibrated experimentally by measuring curvature as a function of strain mismatch in a controlled bilayer beam experiment. However, under the non-uniform thermal actuation considered here, a spatially resolved experimental calibration of $\beta$ is not straightforward and is beyond the scope of the present study. Therefore, in this work we adopt the choice $\beta=1/l_{0}$, such that the dimensionless quantity $\beta l_{0}=1$. This choice minimizes sensitivity to mesh discretization while preserving the correct scaling between strain mismatch and curvature. A more detailed experimental calibration of spatially varying $\beta$ fields will be pursued in future work.
2.2.3 Thermal Strain and Temperature Change Distribution
The thermal strain arises from shrinkage induced by temperature change and is described by the classical relation [47]
$\varepsilon_{\rm pre}^{\rm{th}}=\alpha\Delta T.$ (9)
Experimental observations indicate that deformation near the free boundary of the Shrinky Dink sheet is more pronounced than in regions constrained by the PLA kirigami bilayer. This suggests a spatially nonuniform thermal strain field caused by partial constraint of shrinkage near the bilayer region. To capture this effect, we quantify the distance from an edge midpoint $i$ to the closest Shrinky Dink-PLA bilayer boundary as $d_{i}$, and define $d_{\max}$ as the maximum such distance over the domain,